OpenStudy (samigupta8):

At the time t = 0 the velocity of an object is u and at time t its acceleration f = at. Then the velocity at time t will be A) u B) at C) u + at^2 D) u + ½ at^2

OpenStudy (anonymous):

\[\frac{ dv(t) }{ dt }=at \rightarrow v(t)=\int\limits_{?}^{?}atdt=\frac{ 1 }{ 2 }at^2+C\] \[v(0)=u=C \rightarrow v(t)=u+\frac{ 1 }{ 2 }at^2\]

OpenStudy (samigupta8):

how did u get\[dv(t)\div dt=at\]

OpenStudy (anonymous):

Because it says acceleration at time t is "at" and we know acceleration is the derivative of velocity vs time

OpenStudy (samigupta8):

thankuu.....

OpenStudy (anonymous):

welcome